# Convolutional Hierarchical Deep Learning Neural Networks-Tensor Decomposition (C-HiDeNN-TD): a scalable surrogate modeling approach for large-scale physical systems

### v1

@@ 1 Introduction
Many data-driven methods have been proposed to address the bottleneck of classical numerical solvers (4; 5; 6; 7; 8). Data-driven methods typically adopt the so-called offline-online approach. During the offline stage, high-fidelity data are generated from experiments or numerical simulations. Subsequently, various data-driven models can be used to approximate functional mapping from input to output spaces as surrogates (5; 4; 6; 9). Despite the success of these methods on various problems, a significant challenge that impedes their application to real-world engineering problems is the exorbitant cost of offline simulation data generation for large-scale systems(1).
@@ 2.1 C-HiDeNN
Thanks to the universal approximation theorem, multilayer perceptron (MLP) architectures have been successfully applied in deep learning-based solvers as global basis functions(10). However, MLP-based basis functions need special treatment to satisfy the boundary condition(11). In contrast, finite element (FE) basis functions can easily satisfy the Dirichlet boundary condition by leveraging the Kronecker-Delta property(12). Moreover, numerical integration can be straightforward with Gaussian quadrature. In the previous work, we combined the concept of locally-supported FE basis functions and MLP to develop Convolutional-Hierarchical Deep Learning Neural Networks (C-HiDeNN) basis function. As shown in Eq. 1, $\bm{N_{i}}$ is the linear FE basis function at node $i$; $W$ is the kernel function that can be represented by a partially connected MLP(13; 14). This partially connected MLP is controlled by 3 hyperparameters: connectivity $s$, polynomial order $p$, and dilation parameter $a$. C-HiDeNN has been shown to be a flexible way to construct a high-order approximation with arbitrary convergence rates and automatic mesh adaptivity based on a linear FE mesh(13). More details of the C-HiDeNN formulation can be found in Appendix A.
@@ 2.2 C-HiDeNN Tensor decomposition
Data-driven TD has been extensively investigated as a data compression method for higher-order tensors(15; 17). Unlike data-driven TD, a priori C-HiDeNN-TD can compute the surrogates (solution vectors $\bm{U}_{p_{i}}$) directly from PDE by circumventing the exorbitant offline high-resolution data generation process. Moreover, for high-dimensional problems, C-HiDeNN-TD will only have linear growth of the number of grid points compared to exponential growth in classical numerical algorithms. As a result, C-HiDeNN-TD can directly obtain surrogate models for exa-scale problems (number of grid points $>10^{18}$) with significant speed-up, low memory cost, and efficient disk storage. More details about the C-HiDeNN-TD algorithm can be found in Appendix B.
@@ 3.1 Convergence study: Poisson equation
where the reference solution $u^{ref}$ is obtained using the linear finite element analysis with $8000\times 8000$ elements, as shown in Fig. 1(a).
@@
The convergence plots are shown in Fig. 1(b-d), where number of spatial grid points is defined as the size of the original spatial mesh. It can be seen that the C-HiDeNN-TD has better accuracy compared to the FE-TD (FE basis function is used). Furthermore, C-HiDeNN-TD has a higher accuracy as connectivity ($s$) and polynomial order ($p$) of the neural network increase. This is because larger $s$ and $p$ will lead to higher-order smoothness of the basis function. As can be seen from Fig. 1(d), when the number of ranks increases, the energy norm error will decrease quickly for all cases until it reaches a plateau.
@@ A. C-HiDeNN theory
where $\bm{N}_{i}(\xi)$ is polynomial interpolant at node $i$, $W^{i}_{s,a,p,j}(\xi)$ is the convolution patch function at node $j$ defined on the nodal patch domain $A_{s}^{i}$ with three hyper-parameters ($s$: patch size, $a$: dilation parameter, $p$: polynomial order), and $x^{h,e}(\xi)$ is the FEM mapping from the parent coordinate $\xi$ to the physical coordinate $x$ in element $e$. Note that the input of the neural network is changed from the physical coordinate $x$ to the parent coordinate $\xi$ in Fig. 3(a) to facilitate the numerical integration of the weak form of PDE. The convolution patch functions are precomputed from the meshfree theory and used as weights between the second and third hidden layers in Figure 3(b). More details on how the weights are computed can be found in [Park et al., 2023].
Figure 3: C-HiDeNN structure
@@ B. A priori C-HiDeNN-TD
where $\bm{A_{i}}\subset\mathbb{R}^{ML_{i}\times ML_{i}}$ is the coefficient matrix for dimension $i$ ($M$ refers to number of ranks; $L_{i}$ is the number of nodes in the $i$-th dimension); $\bm{U_{i}}\subset\mathbb{R}^{ML_{i}}$ is the solution vector. Eq. 13a and Eq. 13b are solved iteratively until the variation of $\bm{{U}_{x}}$ and $\bm{{U}_{x}}$ is within the tolerence.

### v2

@@ 1 Introduction
Many data-driven methods have been proposed to address the bottleneck of classical numerical solvers (4; 5; 6; 7; 8). Data-driven methods typically adopt the so-called offline-online approach. During the offline stage, high-fidelity data are generated from experiments or numerical simulations. Subsequently, various data-driven models can be used to approximate functional mapping from input to output spaces as surrogates (5; 4; 6; 9). Despite the success of these methods on various problems, it is still challenging to apply these data-driven methods to large-scale problems where offline training data generation can be very expensive(1).
@@ 2.1 C-HiDeNN
Thanks to the universal approximation theorem, multilayer perceptron (MLP) architectures have been successfully applied in deep learning-based solvers as global basis functions(10). However, MLP-based basis functions need special treatment to satisfy the boundary condition(11). In contrast, finite element (FE) basis functions can automatically satisfy the Dirichlet boundary condition due to the Kronecker-Delta property(12). Moreover, numerical integration can be straightforward with Gaussian quadrature. In the previous work, we combined the concept of locally-supported FE basis functions and MLP to develop Convolutional-Hierarchical Deep Learning Neural Networks (C-HiDeNN) basis function. As shown in Eq. 1, $\bm{N_{i}}$ is the linear FE basis function at node $i$; $W$ is the kernel function that can be represented by a partially connected MLP(13; 14). This partially connected MLP is controlled by 3 hyperparameters: connectivity $s$, polynomial order $p$, and dilation parameter $a$. C-HiDeNN has been shown to be a flexible way to construct a high-order approximation with arbitrary convergence rates and automatic mesh adaptivity based on a linear FE mesh(13). More details of the C-HiDeNN formulation can be found in Appendix A.
@@ 2.2 C-HiDeNN Tensor decomposition
Data-driven TD has been extensively investigated as a data compression method for higher-order tensors(15; 17). Unlike data-driven TD, a priori C-HiDeNN-TD can compute the surrogates (solution vectors $\bm{U}_{p_{i}}$) directly from PDE by circumventing the exorbitant offline high-resolution data generation process. Moreover, for high-dimensional problems, C-HiDeNN-TD will only have linear growth of the number of grid points compared to exponential growth in classical numerical algorithms. As a result, C-HiDeNN-TD can directly obtain surrogate models for exa-scale problems (number of grid points $10^{18}$) with significant speed-up, low memory cost, and efficient disk storage. More details about the C-HiDeNN-TD algorithm can be found in Appendix B.
@@ 3.1 Convergence study: Poisson equation
where the reference solution $u^{ref}$ is obtained using the linear finite element analysis with $8000\times 8000$ elements.
@@
The convergence plots are shown in Fig. 1, where number of spatial grid points is defined as the size of the original spatial mesh. In Fig. 1(a), it can be seen that the C-HiDeNN-TD has better accuracy compared to the FE-TD (FE basis function is used). Furthermore, C-HiDeNN-TD has a higher accuracy as connectivity ($s$) and polynomial order ($p$) of the neural network increase. This is because larger $s$ and $p$ will lead to higher-order smoothness of the basis function. As can be seen from Fig. 1(b), when the number of ranks increases, the energy norm error will decrease quickly for all cases until it reaches a plateau.
@@ A. C-HiDeNN theory
where $\bm{N}_{i}(\xi)$ is polynomial interpolant at node $i$, $W^{i}_{s,a,p,j}(\xi)$ is the convolution patch function at node $j$ defined on the nodal patch domain $A_{s}^{i}$ with three hyper-parameters ($s$: patch size, $a$: dilation parameter, $p$: polynomial order), and $x^{h,e}(\xi)$ is the FEM mapping from the parent coordinate $\xi$ to the physical coordinate $x$ in element $e$. Note that the input of the neural network is changed from the physical coordinate $x$ to the parent coordinate $\xi$ in Fig. 3(a) to facilitate the numerical integration of the weak form of PDE. The convolution patch functions are precomputed from the meshfree theory and used as weights between the second and third hidden layers in Figure 3(b). More details on how the weights are calculated can be found in [Park et al., 2023].
Figure 3: C-HiDeNN structure. (a) 1D C-HiDeNN basis function is constructed by aggregating the information from all the segments patch domain into the current segment (b) Neural network structure of C-HiDeNN basis function: the first 2 hidden layers are mapping the input to piece-wise continuous function NN, which satisfies the Kronecker delta property; the 3rd layer is the nonlinear part that acts as the convolution patch function $W$, which satisfies the reproducing condition; the output field variable $u^{h,e}$ has higher-order smoothness and automatically satisfy the Dirichlet boundary condition thanks to the Kronecker delta property.
@@ B. A priori C-HiDeNN-TD
where $\bm{A_{i}}\subset\mathbb{R}^{ML_{i}\times ML_{i}}$ is the coefficient matrix for dimension $i$ ($M$ refers to number of ranks; $L_{i}$ is the number of nodes in the $i$-th dimension); $\bm{U_{i}}\subset\mathbb{R}^{ML_{i}}$ is the solution vector. Eq. 13a and Eq. 13b are solved iteratively until the variation of $\bm{{U}_{x}}$ and $\bm{{U}_{x}}$ is within the tolerence range.